The line 2x-y+1=0 is tangent to the circle at the point (2,5) and the centre of the circle lies on x-2y=4 then what is the radius of circle??

Grade:12th pass

1 Answers

Rakesh
6 years ago
2x-y+1=0 is a tangent . Slope=-1/2Then eq. Of OA is x+2y=12According to centre of circle of coordinates(x-2y=4)We get,(8,2)R=OA=โˆš(8-2)^2+(2-5)^2=3โˆš5